OpenStudy (anonymous):

What is the value of the function y = 0.5x – 10 when x = 9? y = will fan and medal

OpenStudy (dtan5457):

y=0.5x-10 plug 9 into x y=0.5(9)-10 y=?

OpenStudy (anonymous):

4?

OpenStudy (dtan5457):

0.5 times 9=4.5 y=4.5-10 try again

OpenStudy (anonymous):

um... 3.5? im not sure

OpenStudy (dtan5457):

Calculator anything.. 4.5-10 it has to be a negative number lol

OpenStudy (lizz123):

(hint) use a calculator

OpenStudy (anonymous):

-5.5

OpenStudy (dtan5457):

yes.